OpenStudy (anonymous):

what is the reason behind why we use base 10 rather than base 5,6 etc in calculations

OpenStudy (anonymous):

u can use anything... bt the log table in ur book only has log values of base 10

OpenStudy (unklerhaukus):

we use base ten because we have ten fingers

OpenStudy (anonymous):

i dont agree with @UnkleRhaukus

OpenStudy (anonymous):

the real actaual base i consider is e .. the best one ..

OpenStudy (unklerhaukus):

can you count in base e?

OpenStudy (anonymous):

do u count in 10

OpenStudy (unklerhaukus):

0, 1, 2, 3, 4, 5, 6, 7, 8, 9

OpenStudy (campbell_st):

actually our number system, the hindu- arabic contains 10 digits, so thats why we have a base 10 number system if your number system has only 5 digits you would have a base 5. So ancient cultures had a base 60 system as it allowed them to trade with others who had a different number system

OpenStudy (anonymous):

leave it .. its worth less topic to debate about..

OpenStudy (unklerhaukus):

nonsense @prawal

OpenStudy (anonymous):

there is no necessasity to use 10 u can use anytihing..

OpenStudy (unklerhaukus):

lets try adding some number in base five lets add 15 and 9 in base five this is 30 + 14 =44 converting back to base 10 5x4 + 1x4 20+4 =24

OpenStudy (unklerhaukus):

that was easier than i thought

OpenStudy (anonymous):

so???

OpenStudy (unklerhaukus):

so MATHS

OpenStudy (unklerhaukus):

now im gonna try adding one-hundred-and-five + forty-four 410 + 134 = 1044 \[= 5^3\times1+5^1\times4+5^0\times4\] \[=125+20+4\] \[=145\]

OpenStudy (anonymous):

i did nt came to know about how u got 145

OpenStudy (unklerhaukus):

i mean 149

OpenStudy (unklerhaukus):

lolz

OpenStudy (anonymous):

k even that how u cal, i came to know abt 1044 but not this

OpenStudy (unklerhaukus):

in base 5 5=10

OpenStudy (anonymous):

ok thank you t dnt we try for other base

OpenStudy (unklerhaukus):

ok how about base eleven the numbers \[0, 1,2,3,4,5,6,7,8,9,10,11\]become\[0_{11}, 1_{11},2_{11},3_{11},4_{11},5_{11},6_{11},7_{11},8_{11},9_{11},T_{11},10_{11}\]

OpenStudy (anonymous):

how can we add 23+54 in base 11

OpenStudy (unklerhaukus):

lets add these numbers \[10+77+121+200\] \[=T_{11}+70_{11}+100_{11}+172_{11}\]\[=T_{11}+332_{11}\]\[=341_{11}\] \[=11^2\times3+11\times4+1\]\[=121\times3+44+1\]\[=363+45\]\[=408\]

OpenStudy (unklerhaukus):

\[23+54\] \[=21_{11}+4T_{11}\]\[=70_{11}\]\[=7\times11\]\[=77\]

OpenStudy (anonymous):

can u pls elobrate 2111+4T11 =7011 =7×11 =77

OpenStudy (anonymous):

i am nt geting how to add in base 11

OpenStudy (unklerhaukus):

\[=21_{11}+4T_{11}\] \[= \begin{array}{c} 21 \\ 4T&+\\\hline\\ 60+10\\\end{array}\]\[=70\]

OpenStudy (unklerhaukus):

remember in base eleven 10 is eleven, not ten

OpenStudy (unklerhaukus):

should we try a simpler example or is this starting to make sense?

OpenStudy (anonymous):

pls try simpler example

OpenStudy (unklerhaukus):

\[5+5\]\[=5_{11}+5_{11}\]\[=T_{11}\]\[=10\] ____________ \[5+6\]\[=5_{11}+6_{11}\]\[=10_{11}\]\[=11\]

OpenStudy (anonymous):

how did 5+5 became 5+6 in further step

OpenStudy (unklerhaukus):

the post above is ment to be two different examples

OpenStudy (anonymous):

hey once u check those 2 answers

OpenStudy (unklerhaukus):

?

OpenStudy (anonymous):

10 base 11 is 10 or 11

ganeshie8 (ganeshie8):

10 base 11 is 11 in base 10

OpenStudy (unklerhaukus):

ten \(10=T_{11}\) eleven \(11=10_{11}\)

OpenStudy (anonymous):

ok i got u

OpenStudy (unklerhaukus):

if we were using hexadecimal (base 16) the numbers \[\{0,1,2,3,4,5,6,7,8,9,10,11,12,13,14,15,16\}\] could be represented like this \[\{0_{16},1_{16},2_{16},3_{16},4_{16},5_{16},6_{16},7_{16},8_{16},9_{16},A_{16},B_{16},C_{16},D_{16},E_{16},F_{16},10_{16}\}\]

OpenStudy (anonymous):

so we can calculate in the same way in many bases

OpenStudy (anonymous):

thank u